OpenStudy (anonymous):

The cost of coffee is determined by the type of coffee beans that are mixed together. The cost of a local mix of Arabica and Robusta beans is modeled by the equation below, where A = the pounds of Arabica and R = pounds of Robusta: 700A + 1200R = 1,000,000 If a mixture contains 1000 pounds of Arabica beans, how many pounds of Robusta beans should be in the mixture? R = ________________ pounds

OpenStudy (dan815):

700A + 1200R = 1,000,000 :This equation says that for every point of Arabica its 700 dollars and 1200 dollars for a pound of Robusta he says there is some mixture of the pounds of Arabica and Robusta such taht the total is 1,000,000 And Also A+R=1000 : A is the number of pounds of arabica and R is the number of pounds of robusta

OpenStudy (dan815):

700A + 1200R = 1,000,000 A+R=1000 We must solve for R from these 2 equations